Calculate the PH and POH of pure water at 25 °C ([H+]=1.00×10-7M).


1
Expert's answer
2021-09-12T00:42:35-0400

Since we know that "pH+pOH=pK_w=14", we proceed to calculate the pH and then the pOH:


"[H^+]=1.00×10^{-7}\,M
\\pH=-log[H^+]
\\pH=-log(1.00×10^{-7}\,M)
\\pH=-\cancel{log(10}^{-7})=-(-7)=7
\\pOH=14-pH=14-7=7"


In conclusion, the pH and pOH of pure water at 25 °C are both equivalent to 7.0

Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.